OpenStudy (anonymous):

can someone show me how to exactly solve: cos(1/8pi)

OpenStudy (anonymous):

I know the answer is \[1/2\sqrt{2+\sqrt{2}}\]

OpenStudy (anonymous):

cos(1/8pi) = 1/√2 * (√(1+cos(pi/4)) =1/√2 * (√(1+√2/2)

OpenStudy (anonymous):

let (pi) be 180 therefore (1/8)pi=(1/8)180=cos(45/2)=1/2sqrt(2+sqrt2) as the final answer u posted

OpenStudy (anonymous):

@matricked i dont understand why cos(1/8pi) can be written as 1/√2 * (√(1+cos(pi/4)) can you say which basic formulas or assumptions you use to do this?

OpenStudy (anonymous):

of i am not wrong i have used cos2t = 2(cost)^2 -1 hope u got it